# Jonas E. Arias Federal Reserve Bank of Philadelphia Juan Rubio-Ramírez

### v1

@@ 5.1 Small SVAR of the World Oil Market
Let us begin by describing the model specification in Kilian and Murphy (2014). They model the global market for crude oil using a 4-variable SVAR including the percent change in global crude oil production, a measure of global real activity, the real price of crude oil and the change in global crude oil inventories above the ground. The SVAR is specified at monthly frequency and the estimation sample is 1973:M2–2009:M8. The model includes 24 lags, a constant, and seasonal dummies to remove seasonal variation. Kilian and Murphy (2014) adopt a weak Gaussian-inverse Wishart prior distribution (see e.g. Uhlig, 2005) for the reduced-form parameters.
@@
The results are broadly identical to those in Kilian and Murphy (2014). In particular, a negative flow supply shock causes a persistent decline in global economic activity and oil inventories, and a persistent increase in the real price of oil. The response of oil production is persistently negative. A positive flow demand shock is associated with a persistent increase in global economic activity and it causes a persistent increase the real price of oil as well as a positive response of oil production. Oil production increases sluggishly given the imposed elasticity bounds and it peaks at about one-year after the shock before declining to pre-shock levels. Finally, a positive speculative demand shock causes a persistent increase in the real price of oil and large increase in inventories. Global real activity and oil production persistently decline in response to this shock but the effects are small.
@@
Next we compare the computational time of applying our Gibbs sampler relative to the accept-reject algorithm. Table 2 reports the time (in minutes) per 1,000 effective draw (i.e, time/effective sample size) using Algorithm 2 and the accept-reject algorithm. As highlighted above, we use Chan, Matthes and Yu’s variant of the accept-reject algorithm. To compute the time per effective draw using the Gibbs sampler, first we generate 100,000 draws and save one every 100. This yields 1,000 draws that we store along with the computational time. Second, we use the resulting 1,000 draws to approximate the effective sample size and we use it to compute the time per effective draw. Since the object of interest in SVARs analysis is typically multi-dimensional, that is, researchers are interested in vector of impulse responses, we compute the multivariate effective sample size proposed by Vats, Flegal, and Jones (2019) based on the vector of contemporaneous impulse response functions. [^2] Multivariate effective sample size calculations generally require a longer MCMC sequence as dimensionality increases; therefore, we focus on contemporaneous impulse responses. Our results are robust to including impulse responses beyond horizon 0. In fact, in our application the MCMC draws of impulse responses are less correlated at longer horizons, so the reported effective sample size should be viewed as conservative. We estimate the multivariate effective sample size based on a multivariate batch mean estimator as in Vats, Flegal, and Jones (2019), with a batch size of 100.
The column “Benchmark Model” compares the effective draws per minute under Kilian and Murphy’s specification. As it can be seen the results are comparable: The numbers reported in the table imply that in this case both algorithm take the same time: for both the Gibbs and the accept-reject sampler it would take about 3 hours to obtain 1,000 effective sample draws. [^3] The computations were performed using MATLAB on an Intel Xeon Platinum 8488C processor with 16 active cores running at 2.4 GHz on an x86__\__64 architecture. In order to demonstrate the advantages of our proposed Gibbs sampler algorithm, we consider a researcher that intends to add a bound on the elasticity of oil price demand to the flow supply shock. As demonstrated by Caldara, Cavallo, and Iacoviello (2019) this is an important consideration in SVARs models of the oil market. Following their work we add the ranking restriction that the price elasticity of crude oil demand is restricted to be in a tight interval around the point estimate of $-0.08$ reported by Caldara, Cavallo, and Iacoviello (2019); in particular, we restrict this elasticity to the interval $(-0.09,-0.07)$.
@@
Table 3 provides a summary of the variables and the sign restrictions imposed on their contemporaneous impulse responses to 10 structural shocks. Chan, Matthes, and Yu (2025) only consider the first 8 shocks. We have added the last two to assess the performance of our algorithm. These last two shocks are labeled oil price and consumer sentiment shock, respectively. The SVAR is specified at quarterly frequency, it includes a constant and 5 lags, and the estimation sample is 1973:Q2–2019:Q4. We assume a Minnesota prior for the reduced-form parameters and we set the hyper-parameters following Giannone, Lenza, and Primiceri (2015). Importantly, Chan, Matthes, and Yu (2025) use the asymmetric priors defined in Chan (2022) for the reduced-form parameters, hence we will use these same priors when considering the accept-reject approach.
Turning to the identification, Chan, Matthes, and Yu (2025) identify 8 structural shocks (demand, investment, financial, monetary, government spending, labor supply and wage bargaining) by means of sign restrictions on the contemporaneous impulse responses as well as by ranking restrictions. In total, there are 105 sign restrictions imposed in their baseline specification. When considering our two added shocks, the oil price and a consumer sentiment shocks mentioned above, we increase the number of sign restrictions to 129. When using the Gibbs sampler we obtain 1 million draws and save one every 100.
@@
Let us begin by describing the selected impulse responses to a unit standard deviation expansionary demand shock, shown in Figure 3(a). Red lines depict point-wise posterior medians. Shaded areas represent point-wise 68 percent posterior probability bands. The sign of the impact responses of real GDP, the PCE price index, the federal funds rate, and the unemployment rate are restricted. The remaining horizons as well the response of the non-residential investment and the real wage are unrestricted. As can be seen, the demand shock causes a transient increase in output and prices, and a decrease in the unemployment rate. The federal funds rate increases in response to the shock. The restrictive stance of monetary policy eventually lower economic activity as can be seen for example in the decline of non-residential investment. The real wage decreases in the short-run in response to the shock as nominal wage increases are not enough to offset higher prices—possibly due to sluggish nominal wage adjustment.
@@
Turning to the financial shock, shown in Figure 4(a), it is worth to highlight that this shock is identified with the same sign restrictions used to identify the investment shock except for the impact response of the S&P 500, which in the case of the financial shock is assumed to be positive instead of negative. Overall, the impulse responses are similar except that the decline in non-residential investment after 5 quarters is slightly less pronounced in the case of a financial shock consistent with the positive response of asset prices.
The impulse responses to a unit standard deviation contractionary monetary policy shock are depicted in Figure 4(b). This shock causes the federal funds rate to remain above zero for more than 2 years reflecting inertia in the conduct of monetary policy. Real GDP and prices decline persistently and the unemployment rate jumps upon impact before slowly returning to baseline. Non-residential investment drops upon impact and rounds one year-after the shock in line with a less restrictive monetary policy stance. The real wage increases driven by a decrease of the price level. A notable aspect of these responses is that they suggest that monetary policy can operate with shorter-lags relative to the long-and-variable lags wisdom.
@@
The government spending shock is shown in Figure 5(a). An expansionary one unit standard deviation government spending shock leads to an increase of real GDP for about two quarters and to a long-lasting increase in the price level. To conclude, we discuss the impulse responses to the supply-related structural shocks, that is, the technology, labor supply, and wage bargaining shock. A unit standard deviation positive technology shock leads to a protracted increase in real GDP and non-residential investment, see Figure 5(b). The higher level of output is accompanied with a sustained decline in the unemployment rate and a sustained increase in the real wage. The federal funds rate raises marginally indicating that monetary policy remains roughly neutral in response to technology shocks.
@@
The responses to a unit standard deviation positive labor supply shock are shown in Figure 6(a). This shock induces a hump-shaped response of real GDP and leads to persistently lower prices. The responses to a unit standard deviation negative wage bargaining shock are shown in Figure 6(b). The identifying assumptions for this shock are identical to an expansionary labor supply shock except that the the unemployment rate is assumed to decrease upon impact. When a negative wage bargaining shock occurs, workers experience a decline in their nominal wage alongside a decrease in the unemployment rate. The real wage remains unaffected upon impact as the lower wages are offset by the assumed decrease in the price level. Subsequently, the price level remains below zero inducing an increase in the real wage.
@@
Let us begin by comparing the efficiency of the Gibbs sampler algorithm relative to the accept-reject algorithm when replicating the identification scheme in Chan, Matthes and Yu (2025). Figure 7(a) reports the time (in minutes) per 1,000 effective draws using Algorithm 2 applied to the 35-variable SVAR as a function of the number of shocks identified. We first approximate the multivariate effective sample size 1 million draws and saving one every 100, and second we divide the time required to obtain such draws by the multivariate effective sample size and scale the resulting number by 1,000. In order to assess the computational time as a function of the size of the identified set, we first obtain draws by identifying only the demand shock, then we the investment shock, then the add the financial shock, and so on until we identify the eight shocks in Table 3. As it can be seen, the time per 1,000 effective draws remains computationally feasible as the number of sign restrictions increases.
@@
Figure 7(b) repeats the same figure but using the efficient accept-reject version of Algorithm 1 proposed by Chan, Matthes, and Yu (2025). We do not need to thin the draws because the draws are independent, hence we only produce 1,000 draws. In this case, the timing increases dramatically as demonstrated in the figure. [^4] If we were using the Normal-Inverse-Wishart priors the time would deteriorate even further as drawing from the posterior when using asymmetric priors is more efficient. Let us highlight, that while both codes can be further optimized and the time will vary depending on the machine and the number of variables. Even so, once we fix the number of variables the main message of this figure will not be altered: the accept-reject algorithm can deteriorate sharply as the identified set narrows. Figure 7(c) puts the times together to facilitate the visual comparison.
To emphasize the previous point, let us now consider more shocks to the point that eventually accept-reject can become impractical. To this end, we augment the number of shocks identified in Chan, Matthes, and Yu (2025) by adding the oil price and a consumer sentiment shock described in Table 3. Figures 8(a)-8(d) repeat the exercise shown in Figure 7(a)-7(b) for the case of 9 and 10 shocks. [^5] The time of the accept-reject algorithm for shocks 9 and 10 is extrapolated based on 10 draws. As the reader can see, the time does not increase exponentially when we use the ESS approach. In contrast, when we consider 9 shocks under the accept-reject approach the times are now measures in hours and when we consider to 10 shocks the times are measure in days—Figures 8(c) and 8(d) provide the times.
@@
In order to facilitate the comparison easier, Figure 9 overlays both timings to make clear that our algorithm handle settings (in terms of number of variables and shocks) that the traditional accept-reject cannot handle.

### v2

@@ 5.1 Small SVAR of the World Oil Market
Let us begin by describing the model specification in Kilian and Murphy (2014). They model the global market for crude oil using a 4-variable SVAR featuring the percent change in global crude oil production, a measure of global real activity, the real price of crude oil, and the change in global crude oil inventories above ground. The SVAR is specified at monthly frequency and the estimation sample is 1973:M2–2009:M8. The model includes 24 lags, a constant, and seasonal dummies to remove seasonal variation. Kilian and Murphy (2014) adopt a weak Gaussian-inverse Wishart prior distribution (see e.g. Uhlig, 2005) for the reduced-form parameters.
@@
The results are broadly identical to those in Kilian and Murphy (2014). In particular, a negative flow supply shock causes a persistent decline in global economic activity and oil inventories, and a persistent increase in the real price of oil. The response of oil production is persistently negative. A positive flow demand shock is associated with a persistent increase in global economic activity and it causes a persistent increase the real price of oil as well as a positive response of oil production. Oil production increases sluggishly given the imposed elasticity bounds and it peaks at about one year after the shock before declining to pre-shock levels. Finally, a positive speculative demand shock causes a persistent increase in the real price of oil and large increase in inventories. Global real activity and oil production persistently decline in response to this shock but the effects are small.
@@
Next we compare the computational time of applying our Gibbs sampler relative to the accept-reject algorithm. Table 2 reports the time (in minutes) per 1,000 effective draw (i.e., time/effective sample size) using Algorithm 2 and the accept-reject algorithm. As highlighted above, we use Chan, Matthes and Yu’s variant of the accept-reject algorithm. To compute the time per effective draw using the Gibbs sampler, first we generate 100,000 draws and save one every 100. This yields 1,000 draws that we store along with the computational time. Second, we use the resulting 1,000 draws to approximate the effective sample size and we use it to compute the time per effective draw. Since the object of interest in SVARs analysis is typically multi-dimensional, that is, researchers are interested in vector of impulse responses, we compute the multivariate effective sample size proposed by Vats, Flegal, and Jones (2019) based on the vector of contemporaneous impulse response functions. [^2] Multivariate effective sample size calculations generally require a longer MCMC sequence as dimensionality increases; therefore, we focus on contemporaneous impulse responses. Our results are robust to including impulse responses beyond horizon 0. In fact, in our application the MCMC draws of impulse responses are less correlated at longer horizons, so the reported effective sample size should be viewed as conservative. We estimate the multivariate effective sample size based on a multivariate batch mean estimator as in Vats, Flegal, and Jones (2019), with a batch size of 100.
The column “Benchmark Model” compares the effective draws per minute under Kilian and Murphy’s specification. As can be seen the results are comparable: The numbers reported in the table imply that in this case both algorithms take the same time: for both the Gibbs and the accept-reject sampler it would take about 3 hours to obtain 1,000 effective sample draws. [^3] The computations were performed using MATLAB on an Intel Xeon Platinum 8488C processor with 16 active cores running at 2.4 GHz on an x86__\__64 architecture. In order to demonstrate the advantages of our proposed Gibbs sampler algorithm, we consider a researcher that intends to add a bound on the elasticity of oil price demand to the flow supply shock. As demonstrated by Caldara, Cavallo, and Iacoviello (2019) this is an important consideration in SVARs models of the oil market. Following their work we add the ranking restriction that the price elasticity of crude oil demand is restricted to be in a tight interval around the point estimate of $-0.08$ reported by Caldara, Cavallo, and Iacoviello (2019); in particular, we restrict this elasticity to the interval $(-0.09,-0.07)$.
@@
Table 3 provides a summary of the variables and the sign restrictions imposed on their contemporaneous impulse responses to 10 structural shocks. Chan, Matthes, and Yu (2025) consider only the first 8 shocks. We have added the last two to assess the performance of our algorithm. These last two shocks are labeled oil price and consumer sentiment shock, respectively. The SVAR is specified at quarterly frequency, it includes a constant and 5 lags, and the estimation sample is 1973:Q2–2019:Q4. We assume a Minnesota prior for the reduced-form parameters and we set the hyper-parameters following Giannone, Lenza, and Primiceri (2015). Importantly, Chan, Matthes, and Yu (2025) use the asymmetric priors defined in Chan (2022) for the reduced-form parameters, hence we will use these same priors when considering the accept-reject approach. [^4] We thank Christian Matthes for sharing their replication files with us.
Turning to the identification, Chan, Matthes, and Yu (2025) identify 8 structural shocks (demand, investment, financial, monetary, government spending, technology, labor supply and wage bargaining) by means of sign restrictions on the contemporaneous impulse responses as well as by ranking restrictions. In total, there are 105 sign restrictions imposed in their baseline specification. When considering our two added shocks, the oil price and consumer sentiment shocks mentioned above, we increase the number of sign restrictions to 129. When using the Gibbs sampler we obtain 1 million draws and save one every 100.
@@
Let us begin by describing the selected impulse responses to a unit standard deviation expansionary demand shock, shown in Figure 3(a). Red lines depict point-wise posterior medians. Shaded areas represent point-wise 68 percent posterior probability bands. The sign of the impact responses of real GDP, the PCE price index, the federal funds rate, and the unemployment rate are restricted. The remaining horizons as well the response of the non-residential investment and the real wage are unrestricted. As can be seen, the demand shock causes a transient increase in output and prices, and a decrease in the unemployment rate. The federal funds rate increases in response to the shock. The restrictive stance of monetary policy eventually lowers economic activity as can be seen for example in the decline of non-residential investment. The real wage decreases in the short-run in response to the shock as nominal wage increases are not enough to offset higher prices—possibly due to sluggish nominal wage adjustment.
@@
Turning to the financial shock, shown in Figure 4(a), it is worth highlighting that this shock is identified with the same sign restrictions used to identify the investment shock except for the impact response of the S&P 500, which in the case of the financial shock is assumed to be positive instead of negative. Overall, the impulse responses are similar except that the decline in non-residential investment after 5 quarters is slightly less pronounced in the case of a financial shock consistent with the positive response of asset prices.
The impulse responses to a unit standard deviation contractionary monetary policy shock are depicted in Figure 4(b). This shock causes the federal funds rate to remain above zero for more than 2 years, reflecting inertia in the conduct of monetary policy. Real GDP and prices decline persistently and the unemployment rate jumps upon impact before slowly returning to baseline. Non-residential investment drops upon impact and rounds one year after the shock in line with a less restrictive monetary policy stance. The real wage increases, driven by a decrease of the price level. A notable aspect of these responses is that they suggest that monetary policy can operate with shorter-lags relative to the long-and-variable lags wisdom.
@@
The government spending shock is shown in Figure 5(a). An expansionary one unit standard deviation government spending shock leads to an increase of real GDP for about two quarters and to a long-lasting increase in the price level. To conclude, we discuss the impulse responses to the supply-related structural shocks, that is, the technology, labor supply, and wage bargaining shock. A unit standard deviation positive technology shock leads to a protracted increase in real GDP and non-residential investment, see Figure 5(b). The higher level of output is accompanied by a sustained decline in the unemployment rate and a sustained increase in the real wage. The federal funds rate rises marginally, indicating that monetary policy remains roughly neutral in response to technology shocks.
@@
The responses to a unit standard deviation positive labor supply shock are shown in Figure 6(a). This shock induces a hump-shaped response of real GDP and leads to persistently lower prices. The responses to a unit standard deviation negative wage bargaining shock are shown in Figure 6(b). The identifying assumptions for this shock are identical to an expansionary labor supply shock except that the unemployment rate is assumed to decrease upon impact. When a negative wage bargaining shock occurs, workers experience a decline in their nominal wage alongside a decrease in the unemployment rate. The real wage remains unaffected upon impact as the lower wages are offset by the assumed decrease in the price level. Subsequently, the price level remains below zero, inducing an increase in the real wage.
@@
Let us begin by comparing the efficiency of the Gibbs sampler algorithm relative to the accept-reject algorithm when replicating the identification scheme in Chan, Matthes and Yu (2025). Figure 7(a) reports the time (in minutes) per 1,000 effective draws using Algorithm 2 applied to the 35-variable SVAR as a function of the number of shocks identified. We first approximate the multivariate effective sample size 1 million draws and saving one every 100, and second we divide the time required to obtain such draws by the multivariate effective sample size and scale the resulting number by 1,000. In order to assess the computational time as a function of the size of the identified set, we first obtain draws by identifying only the demand shock, then we add the investment shock, then the financial shock, and so on until we identify the eight shocks in Table 3. As can be seen, the time per 1,000 effective draws remains computationally feasible as the number of sign restrictions increases.
@@
Figure 7(b) repeats the same figure but using the efficient accept-reject version of Algorithm 1 proposed by Chan, Matthes, and Yu (2025). We do not need to thin the draws because the draws are independent, hence we produce only 1,000 draws. In this case, the timing increases dramatically as demonstrated in the figure. [^5] If we were using the normal-inverse-Wishart priors the time would deteriorate even further as drawing from the posterior when using asymmetric priors is more efficient. Let us highlight that even though both codes (for the Gibbs sampler and the accept-reject sampler) can be further optimized and the time will vary depending on the hardware architecture and the number of variables, once we fix the number of variables the main message emerging from comparing Figures 7(a)-7(b) will not be altered: the accept-reject algorithm can deteriorate sharply as the identified set narrows. Figure 7(c) puts the times together to facilitate the visual comparison.
To emphasize the previous point, let us now consider more shocks to the point that eventually accept-reject can become impractical. To this end, we augment the number of shocks identified in Chan, Matthes, and Yu (2025) by adding the oil price and a consumer sentiment shock described in Table 3. Figures 8(a)-8(d) repeat the exercise shown in Figure 7(a)-7(b) for the case of 9 and 10 shocks. [^6] The time of the accept-reject algorithm for shocks 9 and 10 is extrapolated based on 10 draws.
@@
As the reader can see, the time does not increase exponentially when we use the ESS approach. In contrast, when we consider 9 shocks under the accept-reject approach the times are now measured in hours and when we consider to 10 shocks the times are measured in days—Figures 8(c) and 8(d) provide the times. In order to facilitate the comparison, Figure 9 overlays both timings to make clear that our algorithm handles settings (in terms of number of variables and shocks) that the traditional accept-reject cannot handle.